The magnetic flux near the axis inside an air core solenoid carrying a current I is π×10^-6/2 Wb. The length the solenoid is 60cm. What will be its magnetic moment ? (Lengt solenoid h of solenoid is large as compared with its cross section)

Grade:12

1 Answers

Manisha Kumari
5 years ago
   B=uoNI/L   flux =B.A   A=flux/B  M=NIA = NI×flux×L/uoNI= flux × L/uo       =π×10^-6×60×10^-2/2×4π×10^-7=0.75 Am^2